# Hierarchical Reasoning Model

### v1

@@ Abstract
22footnotetext: Tsinghua University Corresponding author. Contact: research@sapient.inc
@@ 1 Introduction
The LLMs literature has relied largely on Chain-of-Thought (CoT) prompting for reasoning 10. CoT externalizes reasoning into token-level language by breaking down complex tasks into simpler intermediate steps, sequentially generating text using a shallow model 11. However, CoT for reasoning is a crutch, not a satisfactory solution. It relies on brittle, human-defined decompositions where a single misstep or a misorder of the steps can derail the reasoning process entirely 12, 13. This dependency on explicit linguistic steps tethers reasoning to patterns at the token level. As a result, CoT reasoning often requires significant amount of training data and generates large number of tokens for complex reasoning tasks, resulting in slow response times. A more efficient approach is needed to minimize these data requirements 14.
@@
The human brain provides a compelling blueprint for achieving the effective computational depth that contemporary artificial models lack. It organizes computation hierarchically across cortical regions operating at different timescales, enabling deep, multi-stage reasoning 20, 21, 22. Recurrent feedback loops iteratively refine internal representations, allowing slow, higher-level areas to guide—and fast, lower-level circuits to execute—subordinate processing while preserving global coherence 23, 24, 25. Notably, the brain achieves such depth without incurring the prohibitive credit-assignment costs that typically hamper recurrent networks from backpropagation through time 19, 26.
@@
Furthermore, we proposed a one-step gradient approximation for training HRM, which offers improved efficiency and eliminates the requirement for BPTT. This design maintains a constant memory footprint ($O(1)$ compared to BPTT’s $O(T)$ for $T$ timesteps) throughout the backpropagation process, making it scalable and more biologically plausible.
@@ 2 Hierarchical Reasoning Model
loss = softmax_cross_entropy(y_hat, y)
@@ 3.1 Benchmarks
ARC-AGI Challenge The ARC-AGI benchmark evaluates general fluid intelligence through IQ-test-like puzzles that require inductive reasoning 27. The initial version, ARC-AGI-1, presents challenges as input-output grid pairs that force AI systems to extract and generalize abstract rules from just a few examples. Each task provides a few input–output example pairs (usually 2–3) and a test input. An AI model has two attempts to produce the correct output grid. Although some believe that mastering ARC-AGI would signal true artificial general intelligence, its primary purpose is to expose the current roadblocks in AGI progress. In fact, both conventional deep learning methods and chain-of-thought techniques have faced significant challenges with ARC-AGI-1, primarily because it requires the ability to generalize to entirely new tasks 28.
@@
Sudoku-Extreme Sudoku is a $\times9$ logic puzzle, requiring each row, column, and $\times3$ block to contain the digits 1–9 exactly once. A prediction is determined correct if it exactly matches the puzzle’s unique solution. Sudoku’s complex logical structure makes it a popular benchmark for evaluating logical reasoning in machine learning 62, 63, 64.
The most frequently used Sudoku datasets in research, namely the Kaggle dataset 65, can be fully solved using elementary single-digit techniques 66. The minimal 17-clue puzzles 62, another widely-used collection, might seem more challenging due to its small number of clues. However, this perception is misleading—since 17 represents the minimum number of clues required to guarantee a unique Sudoku solution, these hints need to be highly orthogonal of each other. This orthogonal arrangement leads to many direct, easily-resolved solution paths 67.
@@
We measure puzzle difficulty by counting the number of search backtracks ("guesses") required by a smart Sudoku solver program tdoku, which uses propositional logic to reduce the number of guesses 67. Our Sudoku-Extreme dataset exhibits a mean difficulty of $22$ backtracks per puzzle, significantly higher than existing datasets, including recent handmade puzzles Sudoku-Bench 68 which average just $0.45$ backtracks per puzzle. These subset complexity levels are shown in Figure˜6-(d).
@@ 3.2 Evaluation Details
For all benchmarks, HRM models were initialized with random weights and trained in the sequence-to-sequence setup using the input–output pairs. The two-dimensional input and output grids were flattened and then padded to the maximum sequence length. The resulting performance is shown in Figure˜1. Remarkably, HRM attains these results with just ~1000 training examples per task—and without pretraining or CoT labels.
For ARC-AGI challenge, we start with all input-output example pairs in the training and the evaluation sets. The dataset is augmented by applying translations, rotations, flips, and color permutations to the puzzles. Each task examples is prepended with a learnable special token that represents the puzzle it belongs to. At test time, we proceed as follows for each test input in the evaluation set: (1) Generate and solve 1000 augmented variants and, for each, apply the inverse‐augmentation transform to obtain a prediction. (2) Choose the two most popular predictions as the final outputs. [^3] The ARC-AGI allows two attempts for each test input. All results are reported on the evaluation set.
@@
In Figure˜1, the baselines are grouped based on whether they are pre-trained and use CoT, or neither. The "Direct pred" baseline means using "direct prediction without CoT and pre-training", which retains the exact training setup of HRM but swaps in a Transformer architecture. Interestingly, on ARC-AGI-1, "Direct pred" matches the performance of Liao and Gu 73, who built a carefully designed, domain-specific equivariant network for learning the ARC-AGI task from scratch, without pre-training. By substituting the Transformer architecture with HRM’s hierarchical framework and implementing ACT, we achieve more than a twofold performance improvement.
@@ 3.3 Visualization of intermediate timesteps
Importantly, the model shows it can adapt different reasoning approaches, likely choosing an effective strategy for each particular task. Further research is needed to gain more comprehensive insights into these solution strategies.
@@ 4 Brain Correspondence
A key principle from systems neuroscience is that a brain region’s functional repertoire—its ability to handle diverse and complex tasks—is closely linked to the dimensionality of its neural representations 75, 76. Higher-order cortical areas, responsible for complex reasoning and decision-making, must handle a wide variety of tasks, demanding more flexible and context-dependent processing 77. In dynamical systems, this flexibility is often realized through higher-dimensional state-space trajectories, which allow for a richer repertoire of potential computations 78. This principle gives rise to an observable dimensionality hierarchy, where a region’s position in the processing hierarchy correlates with its effective dimensionality. To quantify this phenomena, we can examine the Participation Ratio (PR) which serves as a standard measure of the effective dimensionality of a high-dimensional representation 79. The PR is calculated using the formula
@@
The dimensionality hierarchy can be observed, for example, in the mouse cortex, where the PR of population activity increases monotonically from low-level sensory areas to high-level associative areas, supporting this link between dimensionality and functional complexity 74 (Figure˜8 a,b).
We evaluated whether HRM reproduces this neuroscientific principle by calculating the PR for both recurrent modules after training on the Sudoku-Extreme Full dataset. The PR computation used the covariance matrix derived from neural states gathered across multiple Sudoku-solving trajectories. The results show a striking parallel to the biological findings. The low-level module’s state ($z_{L}$) occupies a relatively small subspace with a participation ratio of 30.22, whereas the high-level module’s state ($z_{H}$) operates in a substantially larger subspace with a participation ratio of 89.95, as shown in Figure˜8-(c). Furthermore, Figure˜8-(d) shows that increasing the number of unique tasks (trajectories) from 10 to 100 causes $z_{H}$ dimensionality to scale up accordingly, while $z_{L}$ dimensionality remains stable. These results suggest an emergent separation of representational capacity between the modules that parallels their functional roles.
@@
The results, shown in Figure˜8-(e,f), reveal a stark contrast: the high-level and low-level modules of the untrained network exhibit no hierarchical separation, with their PR values remaining low and nearly indistinguishable from each other. This control analysis validates that the dimensionality hierarchy is an emergent property that arises as the model learns to perform complex reasoning.
@@ 5 Related Work
The hierarchical multi-timescale structure plays also an important role in how the brain processes memory. Models such as Hierarchical Sequential Models 93 and Clockwork RNN 94 use multiple recurrent modules that operate at varying time scales to more effectively capture long-range dependencies within sequences, thereby mitigating the forgetting issue in RNNs.
@@ 6 Discussions
Beyond fine-tuning using human-annotated CoT, reinforcement learning (RL) represents another widely adopted training methodology. However, recent evidence suggests that RL primarily unlock existing CoT-like capabilities rather than discovering fundamentally new reasoning mechanisms 96, 97, 98, 99. Additionally, CoT-training with RL is known for its instability and data inefficiency, often requiring extensive exploration and careful reward design. In contrast, HRM takes feedback from dense gradient-based supervision rather than relying on a sparse reward signal. Moreover, HRM operates naturally in a continuous space, which is biologically plausible and avoids allocating same computational resources to each token, even though tokens vary in their reasoning and planning complexity 16.

### v2

@@ Abstract
22footnotetext: Tsinghua University Corresponding author. Contact: research@sapient.inc. Code available at: github.com/sapientinc/HRM
@@ 1 Introduction
The LLMs literature has relied largely on Chain-of-Thought (CoT) prompting for reasoning 10. CoT externalizes reasoning into token-level language by breaking down complex tasks into simpler intermediate steps, sequentially generating text using a shallow model 11. However, CoT for reasoning is a crutch, not a satisfactory solution. It relies on brittle, human-defined decompositions where a single misstep or a misorder of the steps can derail the reasoning process entirely 12, 13. This dependency on explicit linguistic steps tethers reasoning to patterns at the token level. As a result, CoT reasoning often requires significant amount of training data and generates a large number of tokens for complex reasoning tasks, resulting in slow response times. A more efficient approach is needed to minimize these data requirements 14.
@@
The human brain provides a compelling blueprint for achieving the effective computational depth that contemporary artificial models lack. It organizes computation hierarchically across cortical regions operating at different timescales, enabling deep, multi-stage reasoning 20, 21, 22. Recurrent feedback loops iteratively refine internal representations, allowing slow, higher-level areas to guide, and fast, lower-level circuits to execute—subordinate processing while preserving global coherence 23, 24, 25. Notably, the brain achieves such depth without incurring the prohibitive credit-assignment costs that typically hamper recurrent networks from backpropagation through time 19, 26.
@@
Furthermore, we propose a one-step gradient approximation for training HRM, which offers improved efficiency and eliminates the requirement for BPTT. This design maintains a constant memory footprint ($O(1)$ compared to BPTT’s $O(T)$ for $T$ timesteps) throughout the backpropagation process, making it scalable and more biologically plausible.
@@ 2 Hierarchical Reasoning Model
loss = softmax_cross_entropy(y_hat, y_true)
@@ 3.1 Benchmarks
ARC-AGI Challenge The ARC-AGI benchmark evaluates general fluid intelligence through IQ-test-like puzzles that require inductive reasoning 27. The initial version, ARC-AGI-1, presents challenges as input-output grid pairs that force AI systems to extract and generalize abstract rules from just a few examples. Each task provides a few input–output example pairs (usually 2–3) and a test input. An AI model has two attempts to produce the correct output grid. Although some believe that mastering ARC-AGI would signal true artificial general intelligence, its primary purpose is to expose the current roadblocks in AGI progress. In fact, both conventional deep learning methods and CoT techniques have faced significant challenges with ARC-AGI-1, primarily because it requires the ability to generalize to entirely new tasks 28.
@@
Sudoku-Extreme Sudoku is a $\times9$ logic puzzle, requiring each row, column, and $\times3$ block to contain the digits 1–9 exactly once. A prediction is considered correct if it exactly matches the puzzle’s unique solution. Sudoku’s complex logical structure makes it a popular benchmark for evaluating logical reasoning in machine learning 62, 63, 64.
The most frequently used Sudoku dataset in research, namely the Kaggle dataset 65, can be fully solved using elementary single-digit techniques 66. The minimal 17-clue puzzles 62, another widely-used collection, might seem more challenging due to its small number of clues. However, this perception is misleading—since 17 represents the minimum number of clues required to guarantee a unique Sudoku solution, these hints need to be highly orthogonal to each other. This orthogonal arrangement leads to many direct, easily-resolved solution paths 67.
@@
We measure puzzle difficulty by counting the number of search backtracks (“guesses”) required by a smart Sudoku solver program tdoku, which uses propositional logic to reduce the number of guesses 67. Our Sudoku-Extreme dataset exhibits a mean difficulty of $22$ backtracks per puzzle, significantly higher than existing datasets, including recent handmade puzzles Sudoku-Bench 68 which average just $0.45$ backtracks per puzzle. These subset complexity levels are shown in Figure˜6-(d).
@@ 3.2 Evaluation Details
For all benchmarks, HRM models were initialized with random weights and trained in the sequence-to-sequence setup using the input-output pairs. The two-dimensional input and output grids were flattened and then padded to the maximum sequence length. The resulting performance is shown in Figure˜1. Remarkably, HRM attains these results with just ~1000 training examples per task—and without pretraining or CoT labels.
For ARC-AGI challenge, we start with all input-output example pairs in the training and the evaluation sets. The dataset is augmented by applying translations, rotations, flips, and color permutations to the puzzles. Each task example is prepended with a learnable special token that represents the puzzle it belongs to. At test time, we proceed as follows for each test input in the evaluation set: (1) Generate and solve 1000 augmented variants and, for each, apply the inverse‐augmentation transform to obtain a prediction. (2) Choose the two most popular predictions as the final outputs. [^3] The ARC-AGI allows two attempts for each test input. All results are reported on the evaluation set.
@@
In Figure˜1, the baselines are grouped based on whether they are pre-trained and use CoT, or neither. The “Direct pred” baseline means using “direct prediction without CoT and pre-training”, which retains the exact training setup of HRM but swaps in a Transformer architecture. Interestingly, on ARC-AGI-1, “Direct pred” matches the performance of Liao and Gu 73, who built a carefully designed, domain-specific equivariant network for learning the ARC-AGI task from scratch, without pre-training. By substituting the Transformer architecture with HRM’s hierarchical framework and implementing ACT, we achieve more than a twofold performance improvement.
@@ 3.3 Visualization of intermediate timesteps
Importantly, the model shows that it can adapt to different reasoning approaches, likely choosing an effective strategy for each particular task. Further research is needed to gain more comprehensive insights into these solution strategies.
@@ 4 Brain Correspondence
A key principle from systems neuroscience is that a brain region’s functional repertoire—its ability to handle diverse and complex tasks—is closely linked to the dimensionality of its neural representations 75, 76. Higher-order cortical areas, responsible for complex reasoning and decision-making, must handle a wide variety of tasks, demanding more flexible and context-dependent processing 77. In dynamical systems, this flexibility is often realized through higher-dimensional state-space trajectories, which allow for a richer repertoire of potential computations 78. This principle gives rise to an observable dimensionality hierarchy, where a region’s position in the processing hierarchy correlates with its effective dimensionality. To quantify this phenomenon, we can examine the Participation Ratio (PR), which serves as a standard measure of the effective dimensionality of a high-dimensional representation 79. The PR is calculated using the formula
@@
The dimensionality hierarchy can be observed, for example, in the mouse cortex, where the PR of population activity increases monotonically from low-level sensory areas to high-level associative areas, supporting this link between dimensionality and functional complexity 74 (Figure˜8 (a,b)).
We evaluated whether HRM reproduces this neuroscientific principle by calculating the PR for both recurrent modules after training on the Sudoku-Extreme Full dataset. The PR computation used the covariance matrix derived from neural states gathered across multiple Sudoku-solving trajectories. The results show a striking parallel to the biological findings. The low-level module’s state ($z_{L}$) occupies a relatively small subspace with a participation ratio of 30.22, whereas the high-level module’s state ($z_{H}$) operates in a substantially larger subspace with a participation ratio of 89.95, as shown in Figure˜8(c). Furthermore, Figure˜8(d) shows that increasing the number of unique tasks (trajectories) from 10 to 100 causes $z_{H}$ dimensionality to scale up accordingly, while $z_{L}$ dimensionality remains stable. These results suggest an emergent separation of representational capacity between the modules that parallels their functional roles.
@@
The results, shown in Figure˜8(e,f), reveal a stark contrast: the high-level and low-level modules of the untrained network exhibit no hierarchical separation, with their PR values remaining low and nearly indistinguishable from each other. This control analysis validates that the dimensionality hierarchy is an emergent property that arises as the model learns to perform complex reasoning.
@@ 5 Related Work
The hierarchical multi-timescale structure also plays an important role in how the brain processes memory. Models such as Hierarchical Sequential Models 93 and Clockwork RNN 94 use multiple recurrent modules that operate at varying time scales to more effectively capture long-range dependencies within sequences, thereby mitigating the forgetting issue in RNNs.
@@ 6 Discussions
Beyond fine-tuning using human-annotated CoT, reinforcement learning (RL) represents another widely adopted training methodology. However, recent evidence suggests that RL primarily unlocks existing CoT-like capabilities rather than discovering fundamentally new reasoning mechanisms 96, 97, 98, 99. Additionally, CoT-training with RL is known for its instability and data inefficiency, often requiring extensive exploration and careful reward design. In contrast, HRM takes feedback from dense gradient-based supervision rather than relying on a sparse reward signal. Moreover, HRM operates naturally in a continuous space, which is biologically plausible and avoids allocating same computational resources to each token, even though tokens vary in their reasoning and planning complexity 16.